OpenStudy (anonymous):

Solve 7x^2/x^2+24x=x

OpenStudy (campbell_st):

multiply both sides by (x^2 + 24) to remove the denominator \[7x^2 = x( x^2 + 24x) \] expand and then collect like terms \[x^3 + 17x^2 = 0\] you can now factorise the cubic to get the solutions.

OpenStudy (anonymous):

thanks!